OpenStudy (anonymous):

Use the Mean Value Theorem to find the value of ‘c’ for the function f(x)= x^2/x+2 in [1, 2]. thanks :)

OpenStudy (anonymous):

\[f(x)=\frac{x^2}{x+2}\] \[f(2)=1\] \[f(1)=\frac{1}{3}\] \[\frac{1-\frac{1}{3}}{2-1}=\frac{2}{3}\]

OpenStudy (anonymous):

\[f'(x)=\frac{x(x+4)}{(x+2)^2}\] \[\frac{x(x+4)}{(x+2)^2}=\frac{2}{3}\] \[3x(x+4)=2(x+2)^2\]

OpenStudy (anonymous):

now i guess we have to solve this quadratic for x. hope i can to it

OpenStudy (anonymous):

lmaooo :P

OpenStudy (anonymous):

\[3x^2+12x=2x^2+8x+8\] \[x^2+4x=8\] \[(x+2)^2=8+4=12\] \[x+2=\pm\sqrt{12}\] \[x=-2\pm2\sqrt{3}\]

OpenStudy (anonymous):

only answer in our interval is \[-2+2\sqrt{3}\]

OpenStudy (anonymous):

but...

OpenStudy (anonymous):

it says the answer is between 1 and 2 :/ I dont get it :(